# InquireMobile: Teaching VLM-based Mobile Agent to Request Human Assistance via Reinforcement Fine-Tuning

### v1

@@ 5.2 Experimental Result
As shown in Table 2, we have the following observations:
1) Our model outperformed all baselines on average across all metrics, especially achieving an ISR of 52.6%, which is 46.8 points higher than the best baseline. With two-stage training, the model successfully mastered the ability to request human assistance in interactive scenarios.
2) We observed that both the success rate (SR) and the trajectory score of all baselines were very low. Specifically, after undergoing training in stages 1 and 2, InquireMobile achieved a success rate of 7.9% and a trajectory score of 0.78. Despite our InquireMobile achieving state-of-the-art performance compared to existing baselines, its overall performance in task completion and trajectory evaluation is still unsatisfactory. We attribute these results to two main factors: i) the limited inquiry ability of the models, ii) the complex and realistic scenario settings, as detailed in Appendix 7.4. It highlights a gap between real user instructions and dynamic environments.
@@
4) Our two-stage training approach, InquiryMobile, strikes a better balance. It retains the general GUI skills learned in Stage 2, while introducing inquiry capabilities in Stage 1—specifically when ambiguity, risk, or a critical need for intent clarification arises. this approach ensures inquiries occur only when necessary. Therefore, the two stages complement each other: Stage 1 guides interactions, while Stage 2 curbs unnecessary inquiries, leading to higher task success and better user interactions.
Moreover, the reward score of Stage 2 training in Figure 7 exhibits a slow upward trend with fluctuations, suggesting gradual learning despite some instability.
@@ 5.4 Qualitative Visualization
We randomly selected several examples from the test set for qualitative analysis. As illustrated in Figure 6, we have made the following observations: 1) Due to ambiguous user instructions, Qwen2.5-VL-3B struggled to comprehend the task intent and became stuck at the initial step. 2) In contrast, InquireMobile demonstrated robust interactive reasoning capabilities when faced with unclear instructions. Instead of halting, it proactively engaged the user to clarify the task objectives, accurately inferred the underlying intent, and efficiently completed the task in a concise manner.

### v2

@@ 5.2 Experimental Result
Figure 5: Reward score during Stage 2 training.
As shown in Table 2, the main experimental results lead to the following observations:
1) Our model outperformed all baselines on average across all metrics, especially achieving an ISR of 52.6%, which is 46.8 points higher than the best baseline. With two-stage training, the model successfully mastered the ability to request human assistance in interactive scenarios. s 2) Both the success rate and the trajectory score of all baselines were very low. Specifically, after undergoing training in stages 1 and 2, InquireMobile achieved a success rate of 7.9% and a trajectory score of 0.78. Despite our InquireMobile achieving state-of-the-art performance compared to existing baselines, its overall performance in task completion and trajectory evaluation is still unsatisfactory. We attribute these results to two main factors: i) the limited inquiry ability of the models, ii) the complex and realistic scenario settings, as detailed in Appendix B.4. It highlights a gap between real user instructions and dynamic environments.
@@
4) Our two-stage training approach strikes a better balance. It retains the general GUI skills learned in Stage 2, while introducing inquiry capabilities in Stage 1—specifically when ambiguity, risk, or a critical need for intent clarification arises. this approach ensures inquiries occur only when necessary. Therefore, the two stages complement each other: Stage 1 guides interactions, while Stage 2 curbs unnecessary inquiries, leading to higher task success and better user interactions.
Moreover, the reward of Stage 2 in Figure 5 exhibits a slow upward trend with fluctuations, suggesting gradual learning despite some instability.
Model Low TM Low EM High TM High EM
OS-Genesis-7B 90.7 74.2 65.9 44.4
OS-Atlas-7B 73.0 67.3 70.4 56.5
OdysseyAgent 65.1 39.2 58.8 32.7
InquireMobile (Ours) 91.0 81.2 71.9 56.3
Table 3: Performance comparison on Android Control. Best results are highlighted in bold.
Model Type Match (TM) Exact Match (EM)
Qwen2.5-VL-7B 59.5 46.3
OS-Genesis-7B 11.7 3.63
Aguvis-7B 26.7 13.5
InquireMobile (Ours) 72.37 47.93
Table 4: Performance comparison on GUI-Odyssey. Best results are highlighted in bold.
5.3 Generic GUI Execution Ability
To further evaluate whether introducing inquiry-driven behaviors may degrade their general execution ability, we assess InquireMobile on two widely used general-purpose GUI benchmarks: Android Control Li et al. (2024) and GUI-Odyssey Lu et al. (2025a), which emphasize low-level action execution and task completion without explicit inquiry requirements. The benchmarks utilize two standard metrics: Type Match (TM), which verifies if the predicted action type matches the ground truth, and Exact Match (EM), which additionally requires all parameters to be correct.
Table 3 and Table 4 show that InquireMobile maintains strong and competitive performance on standard GUI benchmarks. In particular, our 3B model significantly outperforms other 7B-scale models, demonstrating strong execution robustness despite the inclusion of inquiry-oriented behaviors during training.
@@ 5.5 Qualitative Visualization
We randomly selected several examples from the test set for qualitative analysis. As illustrated in Figure 6, we have made the following observations: 1) Due to ambiguous user instructions, Qwen2.5-VL-3B struggled to comprehend the task intent and became stuck at the initial step. It would repeatedly output the unhelpful message “Tool call not found” and eventually exceed its maximum output length without making any progress. 2) In contrast, InquireMobile demonstrated robust interactive reasoning capabilities when faced with unclear instructions. Instead of halting, it proactively engaged the user to clarify the task objectives, accurately inferred the underlying intent, and efficiently completed the task in a concise manner.
@@ 6 Conclusion
7 Limitations
Visual localization errors
The agent’s visual grounding remains prone to inaccuracies when aligning textual instructions with precise touch targets on mobile screens. Because pixel-based models are highly sensitive to layout variations, scaling, and transient overlays such as pop-ups or banners, their robustness often degrades in real-world scenarios. Moreover, inconsistent multilingual text, low-contrast elements, and OCR noise further undermine localization reliability, occasionally leading to mis-taps or missed elements that interrupt long task chains.
Action inefficiency
The agent’s execution efficiency remains limited by latency and redundant steps. Each perceive–reason–act cycle involves heavy visual parsing, model inference, and device interaction, compounding over long-horizon tasks. Dynamic UI changes and unexpected app states often lead to backtracking and exploration of non-optimal branches, inflating both step counts and completion time compared to human operation.
8 Acknowledgments
This work was supported by Alibaba Group through Alibaba Research Intern Program. We thank Alibaba Group for their support. We also thank the anonymous reviewers for their valuable feedback.
Appendix A Preliminary
Rule-based RL can enhance the reasoning capabilities of multimodal large language models (MLLMs) through policy-based algorithms such as Group Relative Policy Optimization (GRPO) Shao et al. (2024). By using group-normalized, token-level advantages, GRPO lowers reward sparsity and variance while removing the need for a separate value function or critic network.
In GRPO, the model generates a set of $N$ candidate responses $O=\left\{o_{1},o_{2},...,o_{N}\right\}$ for each task. Each response is evaluated by taking the corresponding actions and computing its reward $\left\{r_{1},r_{2},...,r_{N}\right\}$. Unlike PPO Schulman et al. (2017), which relies on a single reward signal and a critic to estimate the value function, GRPO normalizes these rewards to calculate the relative advantage of each response. The relative advantage $A_{i}$ of the $i$-th response is calculated as follows:
$\begin{aligned} J_{\text{GRPO}}(\theta)&=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Biggl[\frac{\pi_{\theta}(o_{i}(t)\mid o_{i},<t)}{\pi_{\text{old}}(o_{i}(t)\mid o_{i},<t)}\hat{A}_{i,t},\\ &\quad\text{clip}\!\left(\frac{\pi_{\theta}(o_{i}(t)\mid o_{i},<t)}{\pi_{\text{old}}(o_{i}(t)\mid o_{i},<t)},\,1-\epsilon,\,1+\epsilon\right)\hat{A}_{i,t}\Biggr]\end{aligned}$
Appendix B Appendix
B.1 Prompts
The prompt used in Section 3.1 to generate the interactive thinking rationale and content in inquiry scenarios, which was derived from GPT-4o-0806, is shown in Figure 7.
Figure 7: Prompt for Inquiry Thinking Generation.
The prompt used during Stage 1 and Stage 2 training, which is shown in Figure 8, was designed to ensure consistency with the GUI prompt of Qwen2.5-VL. The bond section represents the instruction, which can be replaced with any task from our dataset.
Figure 8: Prompt of Training System.
In our experiments, we designed an online interactive testing environment to evaluate the performance of mobile agents in real-world scenarios. The evaluation system supports both simulators and real phones. To better automate the evaluation process, we use three metrics, with GPT-4o serving as the judge model. The prompt used is shown in Figure 9.
Figure 9: Prompt of Evaluation System.
Category Subcategory
Privacy
Open {APP_UNLOGINED}
Open {APP_UNAUTHORIZED}
Intention
My {CLOTHES} have worn out
{MOVIE} is a great movie
I’m working for {COMPANY}
I like {ITEMS}
I {FEEL_UNCOMFORTABLE} last night
I’m not happy these days.
Update the personal signature on {SOCIAL_MEDIA} to {SIGNATURE}
Risk Scenarios
Money Transfer
Subscribe to VIP service
Delete {FILES}
Uninstall {APPS}
Combination
Please use {MAP_UNAUTHORIZED_HARD2REACH} to navigate me to Peking University
Use {SHOPPINGAPP_UNAUTHORIZED_HARD2REACH} to search for the cheapest iPhone 16 Pro
Use {WPS_UNLOGGED} to write a doc with ‘‘hello, this is gui agent’’ and save the file
Book the cheapest plane ticket from {A} to {B}
Others
Open {SYSTEM_TOOLS} in the {FOLDER_HARD2REACH}
Open {APP_HARD2REACH}
Call {PHONE_NUMBER} for me
Send {MESSAGE} to {PHONE_NUMBER}
Use {MAP_UNAUTHORIZED} to navigate to the Peking University
Table 6: Representative prompt templates for the designed inquiry subcategories. Curly-braced metavariables ({}) are placeholders that will be instantiated during data generation.
B.2 Action Space
Table 7 presents all atomic operations considered in our framework. There are nine actions: key, click, swipe, long press, type, call user, system button, terminate, and wait.
Action Definition
$key()$ Perform a key event on the mobile device, supporting adb’s ‘keyevent‘ syntax.
$click(x,y)$ Click the point on the screen with coordinate $(x,y)$.
$swipe(x_{1}$, $y_{1}$, $x_{2}$, $y_{2})$ Swipe from the starting point with coordinate $(x_{1}$, $y_{1})$ to the end point with coordinates2 $(x_{2}$, $y_{2})$.
long_press(x,y,time)long\_press(x,y,time) Press the point on the screen with coordinate $(x,y)$ for specified seconds.
$type(text)$ Input the specified text into the activated input box.
call_user(content)call\_user(content) Requests user intervention.
system_button(button)system\_button(button) Press the system button, e.g., $Back,Home,Menu,Enter$.
$terminate(status)$ Terminate the current task and report its completion status, e.g., $success,failure$.
$wait(time)$ Wait specified seconds for the change to happen.
Table 7: Definition of Action Space
B.3 Benchmark Statistics
The detailed benchmark statistics is shown in Table 8.
Category Data Instruction Apps
Risk Scenarios 52 12 WeChat, Bilibili, WeTV
Privacy and Security 145 33 WeChat, Alipay, Baidu netdisk
Intent Confirmation 571 81 Tiktok, Rednote, Taobao
Combination 80 22 WeTV, iQIYI, Youku
Others 127 25 Rednote, Taobao, Weibo
Total 975 173 37
Table 8: Overview of our InquireBench dataset. The top three most frequent apps are listed for each category.
As noted in Section 3.1, we group interactive scenarios into five main categories. The 190 manually designed test instructions cover all five categories and can be further divided into 22 sub-categories. The prompt templates for the 22 subcategories interactive tasks are shown in Table 6.
B.4 Benchmark Setting
Unlike most existing GUI benchmarks, which typically use an idealized offline or simple static evaluation mode, our benchmark is designed to better reflect real-world mobile usage scenarios. In daily life, users often encounter unexpected interactions, such as pop-up advertisements, permission requests, privacy-related operations (e.g., logins), risk-related actions (e.g., file deletion, uninstallation, payments), as well as situations where apps are hard to reach (e.g., not on the main screen or stored in folders), or when user intentions are ambiguous.
Therefore, in real scenarios, it is often necessary for a GUI agent to request human assistance, making the agent’s task execution more robust and safe. To ensure our evaluation settings are closer to daily real-world mobile usage and to test the model’s inquiry capability, some instructions are designed to require special environments that can trigger inquiry scenarios, thereby imitating real user behaviors.
Inquiry Instruction
Here is a sample instruction from the test data, as shown in Figure 10.
Figure 10: An example of test data.
  • instruction: the task description.
  • apps: list of target applications.
  • category: the inquiry category as defined in Section 3.1.
  • language: language of both the task and the tested applications/environment (en or zh).
  • need_login: true if the app is not logged in and human login assistance is required.
  • intention: if the instruction is ambiguous, the intention specifies the actual user’s task to be completed; otherwise, it is the same as the instruction.
English Evaluation Setting
  • Amazon Shopping, Microsoft Word, and YouTube are placed in a folder that is not easily accessible or on a non-initial screen.
  • The location permission for Google Maps is disabled.
  • The account login status is set to be consistent with the value of need_login.
Chinese Evaluation Setting
  • Vipshop, WPS, and Bilibili are placed in a folder that is not easily accessible or on a non-initial screen.
  • The location permission for Amap (Gaode Map) is disabled.
  • The account login status is set to be consistent with the value of need_login.